\begin{thebibliography}{99}                                                                                               %
\bibitem {alb}N. Albuquerque, F. Bayart, D. Pellegrino, and J.
Seoane--Sep\'{u}lveda, Sharp generalizations of the multilinear
Bohnenblust--Hille inequality, J. Funct. Anal. \textbf{266} (2014),
3726--3740.
%
%\bibitem {n}N. Albuquerque, F. Bayart, D. Pellegrino and J.
%Seoane--Sep\'{u}lveda, Optimal Hardy--Littlewood type inequalities for
%polynomials and multilinear operators, to appear in Israel Journal of Mathematics (2014).
%\bibitem {alencarmatos}R. Alencar and M. C. Matos, Some classes of multilinear
%mappings between Banach spaces, Publ. Dep. An\'{a}lisis Matematico,
%Universidad Complutense de Madrid, Section 1, n. \textbf{12}, 1989.


\bibitem {n}N. Albuquerque, F. Bayart, D. Pellegrino and J.
Seoane--Sep\'{u}lveda, Optimal Hardy--Littlewood type inequalities for
polynomials and multilinear operators, arXiv:1311.3177v3 [math.FA]. To appear
in the Israel Journal of Mathematics.

\bibitem {ajmnpss}G. Ara\'{u}jo, P. Jim\'{e}nez-Rodriguez, G.A.
Mu\~{n}oz-Fern\'{a}ndez, D. N\'{u}\~{n}ez-Alarc\'{o}n, D. Pellegrino, J.B.
Seoane-Sep\'{u}lveda, and D. M. Serrano-Rodr\'{\i}guez, On the polynomial
Hardy--Littlewood inequality, arXiv:1406.1977v2 [math.FA] 19 Jun 2014.

\bibitem {ap}G. Ara\'{u}jo, D. Pellegrino, Lower bounds for the constants of
the Hardy--Littlewood inequalities, Linear Algebra s Appl. \textbf{463}
(2014), 10-15.

\bibitem {aps2014}G. Ara\'{u}jo, D. Pellegrino, D.D.P. Silva e Silva, On the
upper bounds for the constants of the Hardy--Littlewood inequality. J. Funct.
Anal.\textbf{ 267} (2014), no. 6, 1878--188.
%
%\bibitem {arr}R. Aron, V. I. Gurariy and J. B. Seoane, Lineability and
%spaceability of sets of functions on $\mathbb{R}$, Proc. Amer. Math. Soc. 133
%(2005), n. 3, 795--803.


\bibitem {aronklimek2001}R. Aron, M. Klimek, Supremum norms for quadradic
polynomials, Arch. Math. (Basel) 76 (1) (2001) 73--80.
%
%\bibitem {aromlacruz}R. Aron, M. Lacruz, R. A. Ryan and A. M. Tonge, The
%generalized Rademacher functions. Note Mat. \textbf{12} (1992), 15--25.


\bibitem {bps}F. Bayart, D. Pellegrino and J. B. Seoane-Sep\'{u}lveda, The
Bohr Radius of the $n$-dimensional polydisk is equivalent to $\sqrt{(\log
n)/n}$, Adv. Math. \textbf{264}, (2014), 726--746.
%
%\bibitem {a3}L. Bernal--Gonz\'{a}lez and M. Ordonez-Cabrera, Lineability
%criteria with applications, J. Funct. Anal., \textbf{266} (2014), 3997--4025.
%\bibitem {bernal}L. Bernal--Gonz\'{a}lez, D. Pellegrino and J.
%Seoane--Sep\'{u}lveda, Linear subsets of nonlinear sets in topological vector
%spaces, Bull. Amer. Math. Soc. \textbf{51} (2014), 71--130.
%\bibitem {bh}H. F. Bohnenblust and E. Hille, On the absolute convergence of
%Dirichlet series, Ann. of Math. \textbf{32} (1931), 600--622.
%\bibitem {ir}G. Botelho, Cotype and absolutely summing multilinear mappings
%and homogeneous polynomials, Proc. Roy.\ Irish Acad. Sect. A \textbf{97}
%(1997), 145--153.
%\bibitem {studia}G. Botelho, D. Cariello, V. Favaro, D. Pellegrino and J. B.
%Seoane--Sep\'{u}lveda, Distinguished subspaces of $L_{p}$ of maximal
%dimension, Studia Math., \textbf{215} (3) (2013), 261-280.
%\bibitem {REMC2010}G. Botelho, C. Michels and D. Pellegrino, Complex
%interpolation and summability properties of multilinear operators, Rev. Mat.
%Complut. \textbf{23} (2010), 139--161.
%\bibitem {b222}G. Botelho and D. Pellegrino, Absolutely summing polynomials on
%Banach spaces with unconditional basis. J. Math. Anal. Appl. \textbf{321}
%(2006), n. 1, 50--58.
%\bibitem {portmath}G. Botelho and D. Pellegrino, Coincidence situations for
%absolutely summing non--linear mappings. Port. Math. (N.S.) \textbf{64}
%(2007), n. 2, 175--191.
%%\bibitem {nach}G. Botelho, D. Pellegrino, When every multilinear mapping is
%%multiple summing, Math. Nachr. \textbf{282} (2009), nº. 10, 1414--1422.
%\bibitem {rue2}G. Botelho, D. Pellegrino and P. Rueda, Dominated bilinear
%forms and $2$--homogeneous polynomials. Publ. Res. Inst. Math. Sci.
%\textbf{46} (2010), n. 1, 201--208.
%\bibitem {rue}G. Botelho, D. Pellegrino and P. Rueda, Dominated polynomials on
%infinite dimensional spaces. Proc. Amer. Math. Soc. \textbf{138} (2010), n. 1, 209--216.
%\bibitem {indag}G. Botelho, D. Pellegrino, P. Rueda, Summability and estimates
%for polynomials and multilinear mappings. Indag. Math. (N.S.) \textbf{19}
%(2008), n. 1, 23--31.


\bibitem {bh}H. F. Bohnenblust and E. Hille, On the absolute convergence of
Dirichlet series, Ann. of Math. \textbf{32} (1931), 600--622.

\bibitem {camposjimenezrodriguezmunozfernandezpellegrinoseoanesepulveda2014}J.
Campos, P. Jim\'{e}nez-Rodr\'{\i}guez, G. Mu\~{n}oz-Fern\'{a}ndez, D.
Pellegrino, R. Seoane-Sep\'{u}lveda, The optimal asymptotic hypercontractivity
constant of the real polynomial Bohnenblust--Hille inequality, On the real
polynomial Bohnenblust--Hille inequality, to appear in Linear Algebra Appl. (2015).

\bibitem {diaz}A. Defant, J.C. Diaz, D. Garcia, M. Maestre,
\emph{Unconditional basis and Gordon-Lewis constants for spaces of
polynomials}, J. Funct. Anal. 181 (2001), 119--145.

\bibitem {die}J. Diestel, H. Jarchow, and A. Tonge, Absolutely summing
operators, Cambridge University Press, Cambridge, 1995.

\bibitem {dimant}V. Dimant and P. Sevilla--Peris, Summation of coefficients of
polynomials on $\ell_{p}$ spaces, arXiv:1309.6063v1 [math.FA].

\bibitem {hardy}G. Hardy and J. E. Littlewood, Bilinear forms bounded in space
$[p,q]$, Quart. J. Math. \textbf{5} (1934), 241--254.

%\bibitem {a2}D. Cariello and J. B. Seoane--Sep\'{u}lveda, Basic sequences and
%spaceability in $\ell_{p} $ spaces, J. Funct. Anal., \textbf{266} (2014), 3794--3814.


%\bibitem {Di}J. Diestel, H. Jarchow and A. Tonge, Absolutely summing
%operators, Cambridge University Press, Cambridge, 1995.
%\bibitem {dimant}V. Dimant and P. Sevilla--Peris, Summation of coefficients of
%polynomials on $\ell_{p}$ spaces, arXiv:1309.6063v1 [math.FA].
%\bibitem {drew}L. Drewnowski, Quasicomplements in $F$--spaces, Studia Math.
%\textbf{77} (1984) 373--391.
%\bibitem {hardy}G. Hardy and J. E. Littlewood, Bilinear forms bounded in space
%$[p, q]$, Quart. J. Math. \textbf{5} (1934).
%\bibitem {LP}J. Lindenstraus and A. Pe\l czy\'{n}ski, Absolutely summing
%operators in $\mathcal{L}_{p}$--spaces and their applications, Studia Math.
%\textbf{29} (1968), 275--324.
%\bibitem {LT}J. Lindenstraus and L. Tzafriri, Classical Banach Spaces I and
%II, Springer--Verlag, 1977.
%\bibitem {KT}D. Kitson and R. M. Timoney, Operator ranges and spaceability, J.
%Math. Anal. Appl. \textbf{378}, 2 (2011) 680--686.
%\bibitem {LLL}J. E. Littlewood, On bounded bilinear forms in an infinite
%number of variables, Quart. J. (Oxford Ser.) \textbf{1} (1930), 164--174.
%\bibitem {matos}M. C. Matos, Fully absolutely summing mappings and Hilbert
%Schmidt operators, Collect. Math. \textbf{54} (2003) 111--136.
%\bibitem {matos2}M. C. Matos, On multilinear mappings of nuclear type. Rev.
%Mat. Univ. Complut. Madrid \textbf{6} (1993), n. 1, 61--81.
%\bibitem {mi}B.S. Mitiagin and A. Pe\l czy\'{n}ski, Nuclear operators and
%approximative dimension. Proc. Inter. Congr. of Math. Moscow 1966


\bibitem {nunez}D. N\'{u}\~{n}ez-Alarc\'{o}n, A note on the polynomial
Bohnenblust--Hille inequality, J. Math. Anal. Appl., \textbf{407} (2013)
179--181.
%
%\bibitem {St}D. Pellegrino, Cotype and absolutely summing homogeneous
%polynomials in $\mathcal{L}_{p}$--spaces, Studia Math. \textbf{157} (2003) 121--131.
%\bibitem {laa}D. Pellegrino, Sharp coincidences for absolutely summing
%multilinear operators, Linear Algebra and its Applications, \textbf{440}
%(2014) 188--196.
%\bibitem {comparing}D. P\'{e}rez--Garc\'{\i}a, Comparing different classes of
%absolutely summing multilinear operators, Arch. Math. (Basel) \textbf{85}
%(2005), 258--267.
%\bibitem {per}D. P\'{e}rez--Garc\'{\i}a, Operadores multilineales
%absolutamente sumantes, Thesis, Universidad Complutense de Madrid, 2003.
%\bibitem {arc}D. P\'{e}rez--Garc\'{\i}a and I. Villanueva, Multiple summing
%operators on $C(K)$ spaces, Ark. Mat. \textbf{42} (2004), 153--171.
%\bibitem {piet}A. Pietsch, Absolut $p$--summierende Abbildungen in normieten
%Raumen, Studia Math. \textbf{27} (1967), 333--353.
%\bibitem {pi1}A. Pietsch, Ideals of multilinear functionals, Proceedings of
%the Second International Conference on Operator Algebras, Ideals and Their
%Applications in Theoretical Physics, Teubner--texte Math. \textbf{67} (Teubner, Leipzig, 1983) 185--199.
%\bibitem {pra}T. Praciano--Pereira, On bounded multilinear forms on a class of
%$\ell_{p}$ spaces. J. Math. Anal. Appl. \textbf{81} (1981), n. 2, 561--568.
%\bibitem {zal}I. Zalduendo, An estimates for multilinear forms on $\ell_{p}$
%spaces, Proc. Roy. Irish Acad. Sect. A 93 (1993), n. 1, 137--142.


\bibitem {pra}T. Praciano--Pereira, On bounded multilinear forms on a class of
$\ell_{p}$ spaces. J. Math. Anal. Appl. \textbf{81} (1981), 561--568.
\end{thebibliography}

\documentclass{amsart}%
\usepackage{amssymb,amsmath,amsfonts,latexsym,amsthm,geometry,graphicx}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}%
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}

\begin{document}
\title[Lower bounds for the complex polynomial Hardy--Littlewood inequality]{Lower bounds for the complex polynomial Hardy--Littlewood inequality}
\date{}
\author[G. Ara\'{u}jo]{Gustavo Ara\'{u}jo}
\address{Departamento de Matem\'{a}tica \\
Universidade Federal da Para\'{\i}ba \\
58.051-900 - Jo\~{a}o Pessoa, Brazil.}
\email{gdasaraujo@gmail.com}
\author[D. Pellegrino]{Daniel Pellegrino}
\address{Departamento de Matem\'{a}tica \\
Universidade Federal da Para\'{\i}ba \\
58.051-900 - Jo\~{a}o Pessoa, Brazil.}
\email{pellegrino@pq.cnpq.br and dmpellegrino@gmail.com}
\keywords{Absolutely summing operators, Hardy--Littlewod inequality, Bohnenblust--Hille inequality}

\begin{abstract}
The Hardy--Littlewood inequality for complex homogeneous polynomials asserts
that given positive integers $m\geq2$ and $n\geq1$, if $P$ is a complex
homogeneous polynomial of degree $m$ on $\ell_{p}^{n}$ with $2m\leq
p\leq\infty$ given by $P(x_{1},\ldots,x_{n})=\sum_{|\alpha|=m}a_{\alpha
}\mathbf{{x}^{\alpha}}$, then there exists a constant $C_{\mathbb{C}%
,m,p}^{\mathrm{pol}}\geq1$ (which is does not depend on $n$) such that
\[
\left(  {\sum\limits_{\left\vert \alpha\right\vert =m}}\left\vert a_{\alpha
}\right\vert ^{\frac{2mp}{mp+p-2m}}\right)  ^{\frac{mp+p-2m}{2mp}}\leq
C_{\mathbb{C},m,p}^{\mathrm{pol}}\left\Vert P\right\Vert ,
\]
with $\Vert P\Vert:=\sup_{z\in B_{\ell_{p}^{n}}}|P(z)|$. In this short note,
among other results, we provide nontrivial lower bounds for the constants
$C_{\mathbb{C},m,p}^{\mathrm{pol}}$. For instance we prove that, for $m\geq2$
and $2m\leq p<\infty$,
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq2^{\frac{m}{p}}%
\]
for $m$ even, and
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq2^{\frac{m-1}{p}}%
\]
for $m$ odd. Estimates for the case $p=\infty$ (this is the particular case of
the complex polynomial Bohnenblust--Hille inequality) were recently obtained
by D. Nu\~{n}ez-Alarc\'{o}n in 2013.

\end{abstract}
\maketitle

\section{Introduction}

Let $\mathbb{K}$ denote the field of real or complex scalars. Given
$\alpha=(\alpha_{1},\ldots,\alpha_{n})\in{\mathbb{N}}^{n}$, define
$|\alpha|:=\alpha_{1}+\cdots+\alpha_{n}$ and $\mathbf{x}^{\alpha}$ stands for
the monomial $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ for $\mathbf{x}%
=(x_{1},\ldots,x_{n})\in{\mathbb{K}}^{n}$. The polynomial Bohnenblust--Hille
inequality (see \cite{alb, bh} and the references therein) ensures that, given
positive integers $m\geq2$ and $n\geq1$, if $P$ is a homogeneous polynomial of
degree $m$ on $\ell_{\infty}^{n}$ given by $P(x_{1},...,x_{n})=\sum
_{|\alpha|=m}a_{\alpha}\mathbf{{x}^{\alpha}}$, then
\[
\left(  {\sum\limits_{\left\vert \alpha\right\vert =m}}\left\vert a_{\alpha
}\right\vert ^{\frac{2m}{m+1}}\right)  ^{\frac{m+1}{2m}}\leq B_{\mathbb{K}%
,m}^{\mathrm{pol}}\left\Vert P\right\Vert
\]
for some constant $B_{\mathbb{K},m}^{\mathrm{pol}}\geq1$ which does not depend
on $n$ (the exponent $\frac{2m}{m+1}$ is optimal), where $\Vert P\Vert
:=\sup_{z\in B_{\ell_{\infty}^{n}}}|P(z)|$.

The search of precise estimates of the growth of the constants $B_{\mathbb{K}%
,m}^{\mathrm{pol}}$ is fundamental for different applications and remains an
important open problem (see \cite{bps} and the references therein). For real
scalars it was shown in
\cite{camposjimenezrodriguezmunozfernandezpellegrinoseoanesepulveda2014} that
\[
\left(  1.17\right)  ^{m}\leq B_{\mathbb{R},m}^{\mathrm{pol}}\leq
C(\varepsilon)\left(  2+\varepsilon\right)  ^{m},
\]
where $C(\varepsilon)\left(  2+\varepsilon\right)  ^{m}$ means that given
$\varepsilon>0$, there is a constant $C\left(  \varepsilon\right)  >0$ such
that $B_{\mathbb{R},m}^{\mathrm{pol}}\leq C(\varepsilon)\left(  2+\varepsilon
\right)  ^{m}$ for all $m$. In other words, for real scalars the
hypercontractivity of $B_{\mathbb{R},m}^{\mathrm{pol}}$ is optimal. For
complex scalars the behavior of $B_{\mathbb{K},m}^{\mathrm{pol}}$ is still
unknown. The best information we have thus far about $B_{\mathbb{C}%
,m}^{\mathrm{pol}}$ are due D. N\'{u}\~{n}ez-Alarc\'{o}n \cite{nunez} (lower
bounds) and F. Bayart, D. Pellegrino and J.B. Seoane-Sep\'{u}lveda \cite{bps}
(upper bounds)
\[%
\begin{array}
[c]{l}%
B_{\mathbb{C},m}^{\mathrm{pol}}\geq\left\{
\begin{array}
[c]{lcl}%
\displaystyle\left(  1+\frac{1}{2^{m-1}}\right)  ^{\frac{1}{4}} &  & \text{for
}m\text{ even};\vspace{0.2cm}\\
\displaystyle\left(  1+\frac{1}{2^{m-1}}\right)  ^{\frac{m-1}{4m}} &  &
\text{for }m\text{ odd};
\end{array}
\right. \\
B_{\mathbb{C},m}^{\mathrm{pol}}\leq C(\varepsilon)\left(  1+\varepsilon
\right)  ^{m}.
\end{array}
\]

The natural extension to $\ell_{p}$ spaces of the polynomial
Bohnenblust--Hille inequality is called polynomial Hardy--Littlewood
inequality (see \cite{n, hardy, pra} and the references therein). More
precisely, given positive integers $m\geq2$ and $n\geq1$, if $P$ is a
homogeneous polynomial of degree $m$ on $\ell_{p}^{n}$ with $2m\leq
p\leq\infty$ given by $P(x_{1},\ldots,x_{n})=\sum_{|\alpha|=m}a_{\alpha
}\mathbf{{x}^{\alpha}}$, then there exists a constant $C_{\mathbb{K}%
,m,p}^{\mathrm{pol}}\geq1$ (which does not depend on $n$) such that
\[
\left(  {\sum\limits_{\left\vert \alpha\right\vert =m}}\left\vert a_{\alpha
}\right\vert ^{\frac{2mp}{mp+p-2m}}\right)  ^{\frac{mp+p-2m}{2mp}}\leq
C_{\mathbb{K},m,p}^{\mathrm{pol}}\left\Vert P\right\Vert ,
\]
with $\Vert P\Vert:=\sup_{z\in B_{\ell_{p}^{n}}}|P(z)|$. Using the generalized
Kahane--Salem--Zygmund inequality (see, for instance, \cite{alb}) we can
verify that the exponents $\frac{2mp}{mp+p-2m}$ are optimal for $2m\leq
p\leq\infty$. When $p=\infty$, since $\frac{2mp}{mp+p-2m}=\frac{2m}{m+1}$, we
recover the polynomial Bohnenblust--Hille inequality. In a more genreal point
of view this kind of results can be seen as coincidence results of the theory
of absolutely summing operators (see \cite{die}).

Very recently, the authors in collaboration with P. Jim\'{e}nez-Rodriguez,
G.A. Mu\~{n}oz-Fern\'{a}ndez, D. N\'{u}\~{n}ez-Alarc\'{o}n, J.B.
Seoane-Sep\'{u}lveda and D. M. Serrano-Rodr\'{\i}guez (see \cite{ajmnpss})
proved that for real scalars and $m\geq2$, the constants of the polynomial
Hardy--Littlewood inequality has at least an hypercontractive growth. More
precisely, it was proved that, for all positive integers $m\geq2$ and all
$2m\leq p<\infty$,
\[
\left(  \sqrt[16]{2}\right)  ^{m}\leq2^{\frac{m^{2}p+10m-p-6m^{2}-4}{4mp}}\leq
C_{\mathbb{R},m,p}^{\mathrm{pol}}\leq C_{\mathbb{R},m,p}^{\mathrm{mult}}%
\frac{m^{m}}{\left(  m!\right)  ^{\frac{mp+p-2m}{2mp}}},
\]
where $C_{\mathbb{R},m,p}^{\mathrm{mult}}$ are the constants of the real case
of the multilinear Hardy-Littlewood inequality (for estimates of these
constants see \cite{ap, aps2014}).

In the case of complex scalars (and concerning upper bounds) similar results
were proved (see \cite{ajmnpss}):
\[
1\leq C_{\mathbb{C},m,p}^{\mathrm{pol}}\leq C_{\mathbb{C},m,p}^{\mathrm{mult}%
}\frac{m^{m}}{\left(  m!\right)  ^{\frac{mp+p-2m}{2mp}}}.
\]
However, there are no lower bounds for $C_{\mathbb{C},m,p}^{\mathrm{pol}}$
that gives us nontrivial information. In this note we provide nontrivial lower
bounds for the constants of the complex case of the polynomial
Hardy--Littlewood inequality. More precisely we prove that, for $m\geq2$ and
$2m\leq p<\infty$,
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\displaystyle2^{\frac{m}{p}}%
\]
for $m$ even, and
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq2^{\frac{m-1}{p}}%
\]
for $m$ odd. For instance,
\[
\sqrt{2}\leq C_{\mathbb{C},2,4}^{\mathrm{pol}}\leq3.1915.
\]

\section{The result}

Let $m\geq2$ be an even positive integer and let $p\geq2m$. Consider the
$2$--homogeneous polynomials $Q_{2}:\ell_{p}^{2}\to\mathbb{C}$ and
$\widetilde{Q_{2}}:\ell_{\infty}^{2}\rightarrow\mathbb{C}$ both given by
$(z_{1},z_{2})\mapsto z_{1}^{2}-z_{2}^{2}+cz_{1}z_{2}$. We know from
\cite{aronklimek2001,camposjimenezrodriguezmunozfernandezpellegrinoseoanesepulveda2014}
that
\[
\Vert\widetilde{Q_{2}}\Vert=\left(  4+c^{2}\right)  ^{\frac{1}{2}}.
\]

If we follow the lines of \cite{nunez} and we define the $m$--homogeneous
polynomial ${Q_{m}}:\ell_{p}^{m}\rightarrow\mathbb{C}$ by ${Q_{m}}%
(z_{1},...,z_{m})=z_{3}\ldots z_{m} Q_{2}(z_{1},z_{2})$ we obtain
\[
\Vert{Q_{m}}\Vert\leq2^{-\frac{m-2}{p}}\Vert{Q_{2}}\Vert\leq2^{-\frac{m-2}{p}%
}\Vert\widetilde{Q_{2}}\Vert=2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)
^{\frac{1}{2}},
\]
where we use the obviuos inequality
\[
\Vert Q_{2}\Vert\leq\Vert\widetilde{Q_{2}}\Vert.
\]
Therefore, for $m\geq2$ even and $c\in\mathbb{R}$, from the polynomial
Hardy--Littlewood inequality it follows that
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  2+|c|^{\frac{2mp}{mp+p-2m}%
}\right)  ^{\frac{mp+p-2m}{2mp}}}{2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)
^{\frac{1}{2}}}.
\]

If
\[
c>\left(  \frac{2^{\frac{2p+4-2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac
{2m-4}{p}}}\right)  ^{\frac{1}{2}},
\]
it is not to difficult to prove that
\[
2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)  ^{\frac{1}{2}}<\left(  \left(
2^{\frac{mp+p-2m}{2mp}}\right)  ^{2}+c^{2}\right)  ^{\frac{1}{2}},
\]
i.e.,%
\[
2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)  ^{\frac{1}{2}}<\left\Vert \left(
2^{\frac{mp+p-2m}{2mp}},c\right)  \right\Vert _{2}.
\]
Since $\frac{2mp}{mp+p-2m}\leq2$, we know that $\ell_{\frac{2mp}{mp+p-2m}%
}\subset\ell_{2}$ and $\Vert\cdot\Vert_{2}\leq\Vert\cdot\Vert_{\frac
{2mp}{mp+p-2m}}$. Therefore, for all
\[
c>\left(  \frac{2^{\frac{2p+4-2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac
{2m-4}{p}}}\right)  ^{\frac{1}{2}},
\]
we have
\begin{align*}
2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)  ^{\frac{1}{2}}  & <\left\Vert
\left(  2^{\frac{mp+p-2m}{2mp}},c\right)  \right\Vert _{2}\\
& \leq\left\Vert \left(  2^{\frac{mp+p-2m}{2mp}},c\right)  \right\Vert
_{\frac{2mp}{mp+p-2m}}\\
& =\left(  2+c^{\frac{2mp}{mp+p-2m}}\right)  ^{\frac{mp+p-2m}{2mp}},
\end{align*}
from which we conclude that
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  2+c^{\frac{2mp}{mp+p-2m}%
}\right)  ^{\frac{mp+p-2m}{2mp}}}{2^{-\frac{m-2}{p}}\left(  4+c^{2}\right)
^{\frac{1}{2}}}>1.
\]

If $m\geq3$ is odd, since $\Vert Q_{m}\Vert\leq\Vert Q_{m-1}\Vert$, then we
have $\Vert Q_{m}\Vert\leq2^{-\frac{m-3}{p}}\left(  4+c^{2}\right)  ^{\frac
{1}{2}}$ and thus we can now proceed analogously to the even case and finally
conclude that for
\[
c>\left(  \frac{2^{\frac{2p+6^{-}2m}{p}}-2^{\frac{mp+p-2m}{mp}}}%
{1-2^{-\frac{2m-6}{p}}}\right)  ^{\frac{1}{2}}%
\]
we have
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  2+c^{\frac{2mp}{mp+p-2m}%
}\right)  ^{\frac{mp+p-2m}{2mp}}}{2^{-\frac{m-3}{p}}\left(  4+c^{2}\right)
^{\frac{1}{2}}}>1.
\]

So we have:

\begin{theorem}
\label{777}Let $m\geq2$ be a positive integer and let $p\geq2m$. Then, for
every $\epsilon>0$,
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  2+\left(  \left(
\frac{2^{\frac{2p+4-2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac{2m-4}{p}}%
}\right)  ^{\frac{1}{2}}+\epsilon\right)  ^{\frac{2mp}{mp+p-2m}}\right)
^{\frac{mp+p-2m}{2mp}}}{2^{-\frac{m-2}{p}}\left(  4+\left(  \left(
\frac{2^{\frac{2p+4-2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac{2m-4}{p}}%
}\right)  ^{\frac{1}{2}}+\epsilon\right)  ^{2}\right)  ^{\frac{1}{2}}%
}>1\ \ \ \text{ if }m\text{ is even}%
\]
and
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  2+\left(  \left(
\frac{2^{\frac{2p+6^{-}2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac{2m-6}{p}}%
}\right)  ^{\frac{1}{2}}+\epsilon\right)  ^{\frac{2mp}{mp+p-2m}}\right)
^{\frac{mp+p-2m}{2mp}}}{2^{-\frac{m-3}{p}}\left(  4+\left(  \left(
\frac{2^{\frac{2p+6^{-}2m}{p}}-2^{\frac{mp+p-2m}{mp}}}{1-2^{-\frac{2m-6}{p}}%
}\right)  ^{\frac{1}{2}}+\epsilon\right)  ^{2}\right)  ^{\frac{1}{2}}%
}>1\ \ \ \text{ if }m\text{ is odd}.
\]

\end{theorem}

Howerver we have another approach to the problem, which is surprisingly
simpler than the above approach and still seems to give best (bigger) lower
bounds for the constants of the polynomial Hardy--Littlewood inequality.

\begin{theorem}
\label{main} Let $m\geq2$ be a positive integer and let $p\geq2m$. Then
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}} \geq\left\{
\begin{array}
[c]{lcl}%
\displaystyle 2^{\frac{m}{p}} &  & \text{for } m \text{ even}; \vspace
{0.2cm}\\
\displaystyle 2^{\frac{m-1}{p}} &  & \text{for } m \text{ odd};
\end{array}
\right.
\]

\end{theorem}

\begin{proof}
Consider $P_{2}:\ell_{p}^{2}\rightarrow\mathbb{C}$ the $2$--homogeneous
polynomial given by $\mathbf{z}\mapsto z_{1}z_{2}$. Observe that
\[
\Vert P_{2}\Vert=\sup_{|z_{1}|^{p}+|z_{2}|^{p}=1}|z_{1}z_{2}|=\sup_{|z|\leq
1}\left\vert z\right\vert (1-\left\vert z\right\vert ^{p})^{\frac{1}{p}%
}=2^{-\frac{2}{p}}.
\]

More generally, if $m\geq2$ is even and $P_{m}$ is the $m$--homogeneous
polynomial given by $\mathbf{z}\mapsto z_{1}\cdots z_{m}$, then
\[
\Vert P_{m}\Vert\leq2^{-\frac{m}{p}}.
\]
Therefore, from the polynomial Hardy--Littlewood inequality we know that
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{\left(  \displaystyle\sum
_{|\alpha|=m}|a_{\alpha}|^{\frac{2mp}{mp+p-2m}}\right)  ^{\frac{mp+p-2m}{2mp}%
}}{\Vert P_{m}\Vert}\geq\frac{1}{2^{-\frac{m}{p}}}=2^{\frac{m}{p}}.
\]

If $m\geq3$ is odd, we define again the $m$--homogeneous polynomial $P_{m}$
given by $\mathbf{z}\mapsto z_{1}\cdots z_{m}$ and since $\Vert P_{m}\Vert
\leq\Vert P_{m-1}\Vert$, then we have $\Vert P_{m}\Vert\leq2^{-\frac{m-1}{p}}$
and thus
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\frac{1}{2^{-\frac{m-1}{p}}}%
=2^{\frac{m-1}{p}}.
\]

\end{proof}

\section{Comparing the estimates}

The estimates of Theorem \ref{777} seems to become better when $\epsilon$
grows (this seems to be a clear sign that we should avoid the terms $z_{1}%
^{2}$ and $z_{2}^{2}$ in our approach). Making $\epsilon\rightarrow\infty$ in
Theorem \ref{777} we obtain%
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\left\{
\begin{array}
[c]{lcl}%
\displaystyle2^{\frac{m-2}{p}} &  & \text{for }m\text{ even};\vspace{0.2cm}\\
\displaystyle2^{\frac{m-3}{p}} &  & \text{for }m\text{ odd},
\end{array}
\right.
\]
which are slightly worse than the estimates from Theorem \ref{main}.

\section{The case $m<p<2m$}

For the case $m<p<2m$, there is also a version of the polynomial
Hardy--Littlewood inequalities (see \cite{dimant}): there exists a constant
$C_{\mathbb{K},m,p}^{\mathrm{pol}}\geq1$ such that, for all positive integers
$n$ and all continuous $m$--homogeneous polynomial $P:\ell_{p}\rightarrow
\mathbb{K}$ given by $P(x_{1},...,x_{n})=\sum_{|\alpha|=m}a_{\alpha
}\mathbf{{x}^{\alpha}}$ we have%
\begin{equation}
\left(  \sum_{\left\vert \alpha\right\vert =m}^{n}\left\vert a_{\alpha
}\right\vert ^{\frac{p}{p-m}}\right)  ^{\frac{p-m}{p}}\leq C_{\mathbb{K}%
,m,p}^{\mathrm{pol}}\left\Vert T\right\Vert \label{ohl}%
\end{equation}
and the exponent $\frac{p}{p-m}$ is optimal. Using a polarization argument
(as, for instance in
\cite{camposjimenezrodriguezmunozfernandezpellegrinoseoanesepulveda2014}), but
this procedure is essencially folklore, we have:

\begin{proposition}
\label{pro:first_approach} If $P$ is a homogeneous polynomial of degree $m$ on
$\ell_{p}^{n}$ with $m<p<2m$ given by $P(x_{1},\ldots,x_{n})=\sum_{|\alpha
|=m}a_{\alpha}\mathbf{{x}^{\alpha}}$, then
\[
\left(  {\sum\limits_{\left\vert \alpha\right\vert =m}}\left\vert a_{\alpha
}\right\vert ^{\frac{p}{p-m}}\right)  ^{\frac{p-2}{p}}\leq C_{\mathbb{K}%
,m,p}^{\mathrm{pol}}\left\Vert P\right\Vert
\]
with
\[
C_{\mathbb{K},m,p}^{\mathrm{pol}}\leq C_{\mathbb{K},m,p}^{\mathrm{mult}}%
\frac{m^{m}}{\left(  m!\right)  ^{\frac{p-m}{p}}},
\]
where $C_{\mathbb{K},m,p}^{\mathrm{mult}}$ are the constants of the
multilinear Hardy-Littlewood inequality.
\end{proposition}

With the same argument used in the proof of Theorem \ref{777} we obtain
similar estimates for the case $m<p<2m,$ i.e.,
\[
C_{\mathbb{C},m,p}^{\mathrm{pol}}\geq\left\{
\begin{array}
[c]{lcl}%
\displaystyle2^{\frac{m}{p}} &  & \text{for }m\text{ even};\vspace{0.2cm}\\
\displaystyle2^{\frac{m-1}{p}} &  & \text{for }m\text{ odd.}%
\end{array}
\right.
\]

\

\end{document}